OpenStudy (anonymous):

Calculate the degree of hardness of a water sample containing 0.001 mole of Mgso4 dissolved in 2 litres of water is ?

OpenStudy (anonymous):

sorry, @open_study1 , I have to revise this chapter.. I forgot the concepts :( but u have to convert to ppm unit

OpenStudy (anonymous):

but how?? the answer should be 50ppm............

OpenStudy (anonymous):

why did you invite me to this thread @open_study1

OpenStudy (anonymous):

yes

OpenStudy (apoorvk):

This is the same old question, asked by you aka @shameer1

OpenStudy (anonymous):

@apoorvk plzz help

OpenStudy (anonymous):

can anyone solve this plzzz

OpenStudy (anonymous):

@kryten

OpenStudy (callisto):

Perhaps.... mass of MgSO4 = 0.001 x (24.3 + 32.1 +16.0x4) = xxx ppm = (mg/L) = (xxx/1000) / 2 = ...

OpenStudy (anonymous):

@Vincent-Lyon.Fr plzz help

OpenStudy (vincent-lyon.fr):

Sorry, I do not know how "degree of hardness" is defined.

OpenStudy (callisto):

Is it like ... ppm?

OpenStudy (vincent-lyon.fr):

ppm of what?

OpenStudy (anonymous):

caso4 i think

OpenStudy (callisto):

Mg2+ in this case? Since Mg2+ in water is considered as hard water... Not sure though :|

OpenStudy (anonymous):

we always have to take cas04...

OpenStudy (anonymous):

sorry caco3

OpenStudy (callisto):

From what I've learnt, both Mg2+ and Ca2+ can make hard water. Moreover, in your case, you're only given MgSO4, how come you can have Ca2+ ?

OpenStudy (anonymous):

yes u r correct but after solving ur eq i got a wrong answer .........

OpenStudy (anonymous):

maybe i have gone wrong

OpenStudy (anonymous):

can u plzz check it

OpenStudy (apoorvk):

yes this is defined as 'ppm'

OpenStudy (callisto):

Instead, do you mind showing your workings? I don't mind helping you check your answer :)

OpenStudy (anonymous):

no of moles =0.0005 mgso4 = 0.0005moles caco3

OpenStudy (callisto):

Where did you get ''no of moles =0.0005 mgso4'' ?

OpenStudy (apoorvk):

Where from did you get "0.0005"?

OpenStudy (apoorvk):

lol

OpenStudy (vincent-lyon.fr):

"yes u r correct but after solving ur eq i got a wrong answer ........." Shameer, this is going too far once again! Code of Conduct of OS states that you have to show your work in order to receive help. Not the other way round. Showing your work does not mean writing down the answers from the answer sheet , but actually explaining with WORDS what you are doing.

OpenStudy (anonymous):

0.0005 moles in 1 litre

OpenStudy (anonymous):

in the question given 2 litre so for 1 litre

OpenStudy (anonymous):

ppm should be calculated in 1litre

OpenStudy (apoorvk):

Okay right, 0.0005 moles in one liter - okay. How many moles of water in 1 liter now?

OpenStudy (vincent-lyon.fr):

Dividing 0.001 by 2 is not what I call "showing your work"! ppm, like %, is not calculated in one litre but in any amount.

OpenStudy (anonymous):

how can i show my work.....plz tell

OpenStudy (callisto):

1. type it out 2. write it, then scan it/take a photo of it, after that, upload it

OpenStudy (apoorvk):

Mna just tell me the no. of moles of water in 1000 gms, or 1000ml or 1 liter of it? What's the molar mass of water first of all?

OpenStudy (apoorvk):

*Man

OpenStudy (anonymous):

18

OpenStudy (anonymous):

MM of water h20 = 18

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

no of moles =55

OpenStudy (apoorvk):

yes, sire. in fact 55.55 to be exact. So, you got no. of moles of water in a liter. You got no. of moles of the impurity in a liter too. Can you find the ppm of the impurity in the sample now?

OpenStudy (anonymous):

how?

OpenStudy (apoorvk):

ppm = parts per million. just like percentage is 'parts per 100' helps?

OpenStudy (anonymous):

have to multiply by 10^6???????

OpenStudy (apoorvk):

Yup.. show you working once.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0.0005/55*10^6 is that correct

OpenStudy (apoorvk):

You mean that whole fraction multiplied by a million right?

OpenStudy (anonymous):

s

OpenStudy (apoorvk):

Yup, seems right to me then. @Callisto what do you think?

OpenStudy (anonymous):

no i am not getting the answer the answer should be 50ppm

OpenStudy (apoorvk):

Okay try to calculate everything in terms of gram once - I think that should work. Yes, instead of moles, take the grams of CaSO4 present. What do you get?

OpenStudy (apoorvk):

Hmm, seems I was wrong - ppm can mean by weight as well. Sorry for that @open_study1

OpenStudy (apoorvk):

And in the case of degree of hardness, ppm by weight is considered - though it doesn't make sense to me why.

OpenStudy (anonymous):

Mass of MgSo4=120*10^-3 g.....1 Mass of water = 2000g......2 ppm of mgso4 in water = 1*1000000/2=60ppm i think ...

OpenStudy (anonymous):

@SMISHRA can u ans it

OpenStudy (anonymous):

The answer should be 50 ppm

OpenStudy (callisto):

Here it is... no. of mole of MgSO4 in 1L = 0.001/2 = 0.0005 0.0005 mole of MgSO4 = 0.0005 mole of CaCO3 = 0.0005mole of CaCO3 in 1L = 0.0005 (40.1 + 12.0+16.0x3) = 50mg in 1L therefore degree of hardness of water sample = 50ppm

OpenStudy (anonymous):

thanzzzzz

OpenStudy (callisto):

welcome :|

OpenStudy (apoorvk):

mg/L -> so ppm W/V right? But ppm is supposed to be dimensionless - then how this? o.O I am utterly confused right now :/

OpenStudy (apoorvk):

thanks a lot @Callisto - that cleared all my misconceptions - I was grossly wrong about this. :/

OpenStudy (callisto):

Welcome :)